When atoms are placed at the corners of all 12 edges of a cube how many atoms are present per unit cell?

Grade:12

2 Answers

Arun
5 years ago
Dear student
 
Atoms are present at corner.
no. of atoms = 8 * 1/8 = 1
 
it has 12 atoms at its edge and each contributes 1/4th so 12 * ¼ = 3
 
edge centres cubic unit cell has 4  atoms per unit cell
parul
one year ago
Dear student
The answer must be equal to one
As according to the question, the atoms are present at the corner of 12 edges 
Thus, No. Of atoms in unit cell= 8×1/8 =1
Note-  in question only mentioned that atoms are present at corners but not on edges thus answer will be equal to one